Suppose you deposit $2,368.00 into an account today. In 7.00 years the account is worth $3,988.00. The account earned% per year. Answer Format: Percentage Round to: 2 decimal places (Example: 9.24%, % sign required. Will accept decimal format rounded to 4 decimal places (ex: 0.0924)) Enter Answer Here... Submit Answer
Answer #1

We use the formula

A=P(1+r/100)^n

Where A=future value

P=present value

R is rate of interest

N=time period

3988=2368*(1+r/100)^7

(3988/2368)^(1/7)=1+r/100

1+r/100=1.0773

R=(1.0773-1)*100

=7.73%(approx).